Consider the uniform electric field E =(4.0 ​j​^​​+3.0 ​k​^​​)×10​3^​​ N/C. What is its electric flux through a circular area of radius 1.34 m that lies in the xy-plane? Answer in Nxm^2/C
Answer #1

Electric flux through an area vec{S} due to an electric field vec{E} is Phi= vec{E}cdotvec{S}

Here electric field Ε = (4.0j + 3.0k) * 103 Ус and area vector is rR2 k = (π * 1.342 ) k

Φ-E . S = (4.0j-30k) * 103 . (π * 1.342)A

Φ 3.0 * 10° * π * 1.34-= 1.69 * 104 N. m12/C